What is the solution to the following question "find three consecutive integers whose sum is 303" i missed cass one day and now i am confused with the steps.

Answer by SHUgrad05(58)

Three consecutive integers:
n
n+1
n+2
So:
n+n+1+n+2=303
3n+3=303
3n=300
n=100
n=100; n+1=101; n+2=102